Exercise 8-20

Vaughn Company has used the dollar-value LIFO method for inventory cost determination for many years. The following data were extracted from Vaughn’s records.

Date

Price
Index

Ending Inventory
at Base Prices

Ending Inventory
at Dollar-Value LIFO

December 31, 2017 105 $92,520 $87,250
December 31, 2018 ? 97,520 93,300


Calculate the index used for 2018 that yielded the above results.

Index used for 2018
Answer #1
97520-92520 = 5000
93300-87250 = 6050
Index used for 2018 = 6050/5000*100= 121
Index used for 2018 = 121